OpenStudy (anonymous):

((2x-7)/(x+4))=3 Solve for x

OpenStudy (anonymous):

Will someone please help me I have no idea where to begin

Parth (parthkohli):

Multiply \(x + 4\) to both sides.

OpenStudy (anonymous):

x=11 correct

OpenStudy (anonymous):

Can someone please tell me if my answer is correct

OpenStudy (anonymous):

multiply x+4 over to get \[2x-7 = 3x+12\] isolate x \[3x-2x = -7-12\] \[x = -19\]

OpenStudy (anonymous):

The 2x-7 and x+4 is a fraction they are over each other and =3

OpenStudy (anonymous):

Look at what I did again

OpenStudy (anonymous):

Oh I gotcha I did my math wrong it does come out to -19 for the answer